OpenStudy (anonymous):

What is the value of f[g(3)] for the functions f(x) = x + 3 and g(x) = 2x + 1?

OpenStudy (anonymous):

Any help. im so lost

OpenStudy (anonymous):

Can you find \(g(3)\)?

OpenStudy (anonymous):

I think what you do is substitute 3 in g(x), so g(3) = 2(3) +1 = 7 and then substitute this number in f(g(x)), which would be f(7) = 7+3=10

OpenStudy (anonymous):

then how would you find the vlue

OpenStudy (anonymous):

or would the value be 10?

OpenStudy (anonymous):

so what will my final answer be?

OpenStudy (anonymous):

Yes, 10 would be your final answer.

OpenStudy (anonymous):

alright thank you

OpenStudy (anonymous):

You're welcome!

OpenStudy (anonymous):

Did you understand what I did?

OpenStudy (anonymous):

yes i did. i just couldnt understandhow to start it

OpenStudy (anonymous):

Ok!

OpenStudy (anonymous):

can you help with another question tho

OpenStudy (anonymous):

Let me see, maybe I can help.

OpenStudy (anonymous):

What is the slope of the line through the points (-2, 7) and (3, -3)? m = _?_

OpenStudy (anonymous):

Well, the slope is only the change in Y divided by the change in X. So, M= (y2-y1)/(x2-x1)=(3-(-2))/((-3-7)) = 5/-10 = -1/2

OpenStudy (anonymous):

oh alright thank you that works it out easier

OpenStudy (anonymous):

that makes more sense

OpenStudy (anonymous):

and can you help with this problem What is the value of f(6) for the function f(x) = 3x – 5?

OpenStudy (anonymous):

Well, they are telling you that for any value of x in the function, you will substitute 6 in it. so that means that f(6) = 3(6)-5=13

OpenStudy (anonymous):

oh thaT make so much more sense so my answer is 13 right?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

then how would you do a problem like this What is the inverse of the equation y = 5x - 1

OpenStudy (anonymous):

You'd change the x for the y. You would get X=5y-1 and solve for y. so 5y-x=1 5y=1+x y=1+x/5

OpenStudy (anonymous):

alright i was just making sure thanks

OpenStudy (anonymous):

Alright then!